OpenStudy (anonymous):

I want to know where the volume of the sphere came from. Can some one explain this?

OpenStudy (anonymous):

I know of a Calc III-IV level explanation, but I'm not sure how well I could articulate it...

OpenStudy (e.cociuba):

http://www.k6-geometric-shapes.com/volume-of-a-sphere.html I hope this link helps u!

OpenStudy (anonymous):

Not sure if it's CALC 3, but basically you take a circle and revolve it around the x-axis \[\Large 2\pi\int_0^r\left(\sqrt{r^2-x^2}\right)^2dx \]

OpenStudy (anonymous):

@Spacelimbus, jeez, I was thinking double/triple integrals. Just me over-thinking things...

OpenStudy (anonymous):

I think one reason because I remember such things is because some of our teachers didn't allow us to use formula books, so I always forgot if it's (4/3) or (3/4), until I learned to compute it on my own *grins*

OpenStudy (anonymous):

thank you everyone

OpenStudy (e.cociuba):

np:)

OpenStudy (whpalmer4):

Here's another vote for knowing how to derive the formulas you use! Even if you don't regularly do it, doing it a few times will help summon the correct formula when needed.

OpenStudy (anonymous):

here is an 8 minute video that explains it perfectly! enjoy! http://www.youtube.com/watch?v=tqfU9mC2yFU